# What Do Latent Action Models Actually Learn?

### v1

@@ 4.4 Analysis 4: Improvements via Auxiliary Action Prediction
This section analyzes the setting when a small dataset of action labeled data $\mathcal{D}_{a}$ is available during training of the LAM. This can be used to help guide the latents to represent controllable changes rather than exogenous noise. Specifically we consider the a simple auxiliary loss, with latents as input, predicting the action labels when available. Nikulin et al. (2025) also provide empirically evidence to show that this is a promising strategy to avoiding focus on ‘distractors’ present in the observations.
Action prediction in linear LAM. Consider a linear prediction head $E\in\mathbb{R}^{d_{a}\times d_{z}}$ at the latent bottleneck, $\hat{\boldsymbol{a}}=E\boldsymbol{z}$, and a corresponding action reconstruction loss $\lVert\hat{\boldsymbol{a}}-\boldsymbol{a}\rVert_{2}^{2}$, giving the overall objective, $\mathcal{L}_{a}:=\mathbb{E}_{\mathcal{D}}\left[\lVert\hat{\boldsymbol{o}}^{\prime}-\boldsymbol{o}^{\prime}\rVert_{2}^{2}\right]+\mathbb{E}_{\mathcal{D}_{a}}\left[\lVert\hat{\boldsymbol{a}}-\boldsymbol{a}\rVert_{2}^{2}\right]$ with $\lambda:=|\mathcal{D}_{a}|/|\mathcal{D}|$.
@@
Following the conditions in Proposition 4.3 and assuming $\boldsymbol{\epsilon}^{T}X=0$ and $d_{z}\geq d_{a}$, optimizing $\mathcal{L}_{a}$ biases the encoder parameter perpendicular to the noise. For an artificial case where $\lambda\to+\infty$, we obtain perfect LAM with $D\boldsymbol{\epsilon}=\boldsymbol{0}$ and $B\boldsymbol{z}=\boldsymbol{q}$.
@@
We provide the proof in Appendix C.3. Note that we assume that $\boldsymbol{\epsilon}^{T}X=0$, which means that the signal $\boldsymbol{q}=X\boldsymbol{a}$ and the noise $\boldsymbol{\epsilon}$ are disentangled. This can correspond to the case, e.g., where $\boldsymbol{\epsilon}$ is background disturbance and $\boldsymbol{q}$ is the table-top manipulation (cf. Case 3 in Section 4.1). The conclusion $D\perp\boldsymbol{\epsilon}$ indicates that noise will not enter the latent (noting that $C=-D$ in this case).
@@
Although Proposition 4.3 considers an artificial case with learn $\lambda$, it indicates that auxiliary action prediction can bias linear LAM to denoise. In our simulation, we induce different levels of noise caused by other agents with the same setting as in Figure 3 (right), and present the result in Figure 4 (right). In contrast to previous results where $\sigma_{exo}$ fails the learning of linear LAM, using only 1% of action label in training leads to successful learning. The simulation results indicate that only a small proportion of action label in training can significantly encourage the latent to encode the real actions and denoise.
@@ 5 Conclusion
Latent action models (LAMs) have become popular in the pre-training phase of many embodied AI models. In contrast to the popularity, detailed analysis into the learnability of LAMs is missing. This paper tries to bridge the gap by proposing a linear LAM model that captures the essence of LAM but remains mathematical tractability. By analyzing linear LAM, we have made several fruitful observations, including justifying the design of LAM training loss by showing that it corresponds to PCA and can succeed in recovering the real action, highlighting in the data collection policy’s impact on the learning of LAM, and showing the functionality of data augmentation and auxiliary action prediction. We acknowledge that we should also notice the gap between linear LAM and practical LAM in terms of both the capacity of the FDM and IDM models and the linear dynamics assumed in linear LAM. Nevertheless, we believe linear LAM can be further used as a quick testbed for researchers to develop other tricks for LAM learning.
@@
  • Bottleneck. Both models have an information bottleneck between in the IDM but this is implemented in different ways. Practical LAMs usually adopt vector quantization Van Den Oord et al. (2017), while linear LAM uses a low continuous dimension $d_{z}\ll d_{o}$.
@@ Appendix B Simulation Set-ups
  • Data augmentation. Data augmentation is a trick that may improve the learnability of LAM mentioned in several previous papers (Chen et al., 2024b; Sun et al., 2024). We implement data augmentation by adding a Gaussian noise $\boldsymbol{\kappa}\sim\mathcal{N}(\mathbf{{0},|\boldsymbol{\kappa}|^{2}I})$ for linear LAM. Note that $\mathbb{V}\text{ar}(\boldsymbol{\kappa})=|\boldsymbol{\kappa}|^{2}$. By default data augmentation is turned off, except for Figure 4 (middle and right).
  • Action prediction. Action prediction is another trick proposed in the previous paper (Nikulin et al., 2025). We implement action prediction by predicting the true action label based on the latent with a learnable linear transformation for a small proportion of the data samples. We denote the ratio of the samples that we access their action labels as $\lambda$. We find that setting $\lambda=1\%$ is enough. By default action prediction is turned off, except for Figure 4 (right).
@@ Appendix C Proofs
Regarding the first term, note that $A$ is full rank $d_{o}$ while $BC$ and $BD$ are of rank $d_{z}$. Since $A$ only appears in this term and it is of greater or equal rank than $BD+BD$, it’s optimal value is $A=I-B(C+D)$, setting the first term is zero. Hence we are left with the middle term.
@@
Note that, under the condition of Proposition 4.1, $(A,B,C,D)$ are over-parameterized. When we adopting data augmentation in Proposition 4.3, we obtain $A=I$ and $C+D=0$, which refines how the condition $A=I-B(C+D)$ is satisfied.
@@
Minimizing this function can be achieved by exactly setting $A=I$ and $C=-D$, which zeros the last three terms simultaneously. In this case, the optimization problem again reduces to PCA $\lVert(BD-I)(\boldsymbol{q}+\boldsymbol{\epsilon})\rVert_{2}^{2}$.
Discussion. In the correlated case $\mathbb{E}[\boldsymbol{o}^{T}(\boldsymbol{q}+\boldsymbol{\epsilon})]\neq 0$, when the third term (the cross term) in (17) cannot be ignored, nevertheless there is encouragement to reach $A=I$ and $C=-D$.
@@
Based on the conclusion in Proposition 4.3, optimizing for the first term in $\mathcal{L}_{a}$ results in $A=I$ and $C=-D$. Since $\boldsymbol{\epsilon}^{T}X=0$, we have $\boldsymbol{q}^{T}\boldsymbol{\epsilon}=0$. We consider the case where $\mathcal{D}$ and $\mathcal{D}_{a}$ come from the same distribution and the availability of action labels are uniformly random. Therefore, we can ignore the expectation outside the RHS (for simplicity) and re-write the loss as,
@@
We decompose $X$ as $X=U\Sigma V^{T}$ with $U\in\mathbb{R}^{d_{o}\times d_{a}}$ (which spans the column space of $X$), $\Sigma\in\mathbb{R}^{d_{a}\times d_{a}}$, and $V\in\mathbb{R}^{d_{a}\times d_{a}}$.
We will show that, when $\lambda\to+\infty$, $D=\left[\begin{matrix}U^{T}\\ 0\end{matrix}\right]$, $B=[U,*]$, and $E=[V\Sigma^{-1},*]$ minimizes $\mathcal{L}_{a}$ where * indicates arbitrary entries and * vanishes when $d_{z}=d_{a}$.
First, this solution zeros the last two terms. For the last term, $D\boldsymbol{\epsilon}=\left[\begin{matrix}U^{T}\\ 0\end{matrix}\right]\boldsymbol{\epsilon}=0$ since $\boldsymbol{\epsilon}^{T}X=0$ and $U$ spans the column space of $X$. For the third term, $EDX=[V\Sigma^{-1},*]\left[\begin{matrix}U^{T}\\ 0\end{matrix}\right]U\Sigma V^{T}=I_{d_{a}}$. Then, since we have determined $D$, the second term becomes $\lVert(BD-I)\boldsymbol{\epsilon}\rVert_{2}^{2}=\lVert\boldsymbol{\epsilon}\rVert_{2}^{2}$. We can choose $B=[U,*]$ to zero the first term since $BDX=[U,*]\left[\begin{matrix}U^{T}\\ 0\end{matrix}\right]X=X$. In this way, we obtain the minimal loss $\mathcal{L}^{*}_{a}=\lVert\boldsymbol{\epsilon}\rVert_{2}^{2}$. We can also verify that $B\boldsymbol{z}=B(C\boldsymbol{o}+D\boldsymbol{o}^{\prime})=BD\boldsymbol{q}=UU^{T}\boldsymbol{q}=\boldsymbol{q}$.

### v3

@@ 4.4 Analysis 4: Improvements via Auxiliary Action Prediction
This section analyzes the setting when a small dataset of action labeled data $\mathcal{D}_{a}$ is available during training of the LAM. This can be used to help guide the latents to represent controllable changes rather than exogenous noise. Specifically we consider the a simple auxiliary loss, with latents as input, predicting the action labels when available. nikulin2025distractor also provide empirically evidence to show that this is a promising strategy to avoiding focus on ‘distractors’ present in the observations.
Action prediction in linear LAM. Consider a linear prediction head $E\in\mathbb{R}^{d_{a}\times d_{z}}$ at the latent bottleneck, $\hat{\boldsymbol{a}}=E\boldsymbol{z}$, and a corresponding action reconstruction loss $\lVert\hat{\boldsymbol{a}}-\boldsymbol{a}\rVert_{2}^{2}$, we optimize this objective, $\mathcal{L}_{a}:=\mathbb{E}_{\mathcal{D}}\left[\lVert\hat{\boldsymbol{o}}^{\prime}-\boldsymbol{o}^{\prime}\rVert_{2}^{2}\right]+\lambda\mathbb{E}_{\mathcal{D}_{a}}\left[\lVert\hat{\boldsymbol{a}}-\boldsymbol{a}\rVert_{2}^{2}\right]$ and where $\lambda=|\mathcal{D}_{a}|/|\mathcal{D}|$.
@@
Following the conditions in Proposition 4.3 and assuming $\mathbb{E}[\boldsymbol{q}^{T}\boldsymbol{\epsilon}]=0$ and $d_{z}\geq d_{a}$, optimizing $\mathcal{L}_{a}$ biases the encoder parameter perpendicular to the noise. For an artificial case where $\lambda\to+\infty$, we obtain perfect LAM with $D\boldsymbol{\epsilon}=\boldsymbol{0}$ and $B\boldsymbol{z}=\boldsymbol{q}$.
@@
We provide the proof in Appendix C.3. Note that we assume that $\boldsymbol{\epsilon}^{T}X=0$, meaning that the action effect $\boldsymbol{q}=X\boldsymbol{a}$ and the noise $\boldsymbol{\epsilon}$ are independent. For example, this would hold in the case of table-top manipulation with passers-by occasionally walking by in the background as the source of $\boldsymbol{\epsilon}$ (cf. Case 3 in Section 4.1). The conclusion $D\perp\boldsymbol{\epsilon}$ indicates that noise will not enter the latent (noting that C=DC=-D in this case).
@@
Proposition 4.4 indicates that auxiliary action prediction helps linear LAM to reduce the noise in its latents. Figure 6 (right) present our simulation, we apply different levels of ‘other agents’ noise (case 3 in Section 4.1) with the same settings in Figure 3 (right). Unlike Figure 3 (right), where $\sigma_{exo}=2.0$ caused linear LAM to fail, now 1% of action labels in auxiliary action prediction leads to successful learning. The simulation results indicate that only a small proportion of action labels are needed to encourage the latent to focus on encoding real actions, and not noise.
5 Beyond Linear LAM Experiments
Figure 7: Overview of the practical LAM experiment set up in Section 5. An image of $4\times 4$ pixels is divided into two regions, the top three rows contain controllable effects, while the last row contains exogenous noise. Non-linear neural networks with a VQ bottleneck form the IDM and FDM.
To test whether the theoretical insights from our linear LAM analysis hold in more realistic settings, we ran empirical experiments using a more realistic LAM setup, on a carefully designed synthetic dataset, which allows us to carefully measures proxies for LLO. Figure 7 provides an overview of the setup. We use small $\times4$ images as input (instead of vectors), non-linear CNNs in IDM and FDM (instead of linear layers), and vector quantization (instead of using a reduced linear dimension as a bottleneck). We will show that the main conclusions in our paper still hold in this more complex setting.
Dataset. We designed a $4\times 4$ grid-world style synthetic dataset. The top $3\times 4$ grid of the observation contains a square (intensity=1.0) that can be controlled with five actions (up, down, left, right, and stay still). The bottom $1\times 4$ grid of the observation contains random Bernoulli noise (with prob 0.5). An intensity parameter controls the noise magnitude (none=0.0, low=1.0, high=2.0).
Policy. By default, we use a uniform policy, where each action is equally probable. For one experiment, we also use a correlated policy, where state and action are correlated. With 95% prob, the action moves the square on a fixed snaking pattern through the grid, and with 5% chance a random action is selected.
Model. For the IDM, we use a small CNN to encode $\boldsymbol{o}$ and $\boldsymbol{o}^{\prime}$, followed by a VQ bottleneck with codebook size of 5 outputting the latent. Finally, for the FDM, a separate UNet takes the latent and previous observation $\boldsymbol{o}$ to output the predicted $\hat{\boldsymbol{o}}^{\prime}$. When predicting actions, codes are preassigned to actions, and latents are trained to minimize L2 distance to their true action code. For data augmentation, we shift the $4\times 4$ image left/right for one grid with periodic padding. Models were trained for 16k updates with the Adam optimizer. Unless specified, we use low stochastic noise, no action prediction, no data augmentation, and five codebook vectors.
Evaluation. Not being linear, it is not possible to measure the mutual information between latent and quantities of interest exactly. Instead, since by design the observation separates the controllable (top half) from stochastic (bottom half) changes, we measure the reconstruction loss on the relevant portion of the observation to assess how what information the latent has captured following training. A lower controllable loss means the latent contains more information about the action, and lower stochastic loss means more noise has been captured. Similar to LLO, this is normalized by the variance of the signal. Note there is unfortunately no straightforward way to measure the information about the observation $\boldsymbol{o}$ in the latent for this non-linear case.
Results. We present experiment results containing controllable loss and stochastic loss in Table 1. Firstly, we train vanilla LAM on different noise levels. Results show that as noise is increased from none to high, the action information in the latent decreases (increasing controllable loss). At the same time, the information about the stochastic noise increases (decreasing stochastic loss). This is consistent with our linear LAM theory in Section 4.1, that latents encode whatever leads to most variance. Secondly, we see that switching from a uniform to a correlated policy reduces the information about actions in the latents, while increasing the information about stochastic noise. This is predicted by our linear LAM theory in Section 4.2, that deterministic data collection policy can lead to degenerated performance. Thirdly, we see that data augmentation improves LAM learning. This is consistent with the conclusion in Section 4.3, that data augmentation can help LAM learning. Finally, we show that incorporating 1% of actions labels into the training process improves over vanilla LAM by increasing information about actions, and reducing information about noise. This is consistent with the conclusion in Section 4.4, that predicting true actions improves LAM learning.
Setting Controllable loss ($\downarrow$) Stochastic loss ($\uparrow$)
No noise 0.624 $\pm$ 0.087
Low noise 0.781 $\pm$ 0.079 0.739 $\pm$ 0.047
High noise 1.046 $\pm$ 0.017 0.607 $\pm$ 0.021
Uniform policy 0.781 $\pm$ 0.079 0.739 $\pm$ 0.047
Correlated policy 1.997 $\pm$ 0.022 0.599 $\pm$ 0.036
No data augmentation 0.781 $\pm$ 0.079 0.739 $\pm$ 0.047
Data augmentation 0.415 $\pm$ 0.020 0.898 $\pm$ 0.049
No action prediction 0.781 $\pm$ 0.079 0.739 $\pm$ 0.047
1% action prediction 0.295 $\pm$ 0.011 0.986 $\pm$ 0.002
Table 1: Ablations of performance of practical LAM under different settings. We present the mean and standard errors of controllable loss (measures prediction loss of true action from latent, lower means more action information) and stochastic loss (measures prediction loss of noise from latents, higher means less noise information) over five seeds. Each group corresponds to the analysis of one sub-section in Section 4 and we bold the losses of the better variant in each group.
@@ 6 Conclusion
Latent action models (LAMs) are increasingly used in the pre-training phase of embodied AI models. However, they have so far been intuitively and empirically motivated, without thorough analysis into their learnability. This paper bridges this gap by proposing to study LAMs through a simple model capturing the essence of LAM training while remaining mathematically tractable.
Through our analysis of linear LAM, we made several observations, including showing that under certain assumptions, it corresponds precisely to PCA. In several noise settings, linear LAM leads to recovery of the true action – this supports the use of real-world LAM in low noise settings, as well as the application of preprocessing steps to reduce noise in a dataset. However, we also showed that when the main cause of variation between consecutive observations is not the controllable signal, LAM latents will capture noise rather than action information. We observed a further danger so far undocumented in the literature – a lack of randomization in a data collection policy can harm the latents learned by LAM. Finally, we provided analytical results justifying techniques emerging in the LAM literature to improve the quality of latents learned – data augmentation and auxiliary action prediction.
We acknowledged that several differences exist between our linear LAM and LAMs in practice, and we conducted a further empirical experiment in a controlled non-linear setting that evidence our theoretical conclusions hold in more realistic scenarios.
@@
  • Bottleneck. Both models have an information bottleneck between in the IDM but this is implemented in different ways. Practical LAMs usually adopt vector quantization van2017neural, while linear LAM uses a low continuous dimension $d_{z}\ll d_{o}$.
@@ Appendix B Simulation Set-ups
  • Data augmentation. Data augmentation is a trick that may improve the learnability of LAM mentioned in several previous papers (chen2024igor; sun2024deltadiffusion). We implement data augmentation by adding a Gaussian noise $\boldsymbol{\kappa}\sim\mathcal{N}(\mathbf{{0},|\boldsymbol{\kappa}|^{2}I})$ for linear LAM. Note that $\mathbb{V}\text{ar}(\boldsymbol{\kappa})=|\boldsymbol{\kappa}|^{2}$. By default data augmentation is turned off, except for Figure 6 (middle and right).
  • Action prediction. Action prediction is another trick proposed in the previous paper (nikulin2025distractor). We implement action prediction by predicting the true action label based on the latent with a learnable linear transformation for a small proportion of the data samples. We denote the ratio of the samples that we access their action labels as $\lambda$. We find that setting $\lambda=1\%$ is enough. By default action prediction is turned off, except for Figure 6 (right).
@@ Appendix C Proofs
Regarding the first term, note that $A$ is full rank $d_{o}$ while $BC$ and $BD$ are of rank $d_{z}$. Since $A$ only appears in this term and it is of greater or equal rank than $BD+BD$, it’s optimal value is A=IB(C+D)A=I-B(C+D), setting the first term is zero. Hence we are left with the middle term.
@@
Note that, under the condition of Proposition 4.1, $(A,B,C,D)$ are over-parameterized. When we adopting data augmentation in Proposition 4.3, we obtain $A=I$ and $C+D=0$, which refines how the condition A=IB(C+D)A=I-B(C+D) is satisfied.
@@
Minimizing this function can be achieved by exactly setting $A=I$ and C=DC=-D, which zeros the last three terms simultaneously. In this case, the optimization problem again reduces to PCA $\lVert(BD-I)(\boldsymbol{q}+\boldsymbol{\epsilon})\rVert_{2}^{2}$.
Discussion. In the correlated case $\mathbb{E}[\boldsymbol{o}^{T}(\boldsymbol{q}+\boldsymbol{\epsilon})]\neq 0$, when the third term (the cross term) in (18) cannot be ignored, nevertheless there is encouragement to reach $A=I$ and C=DC=-D.
@@
Based on the conclusion in Proposition 4.3, optimizing for the first term in $\mathcal{L}_{a}$ results in $A=I$ and C=DC=-D. We consider the case where $\mathcal{D}$ and $\mathcal{D}_{a}$ come from the same distribution and the availability of action labels are uniformly random. Therefore, we can ignore the expectation outside the RHS (for simplicity) and re-write the loss as,
@@
The second line leverage the fact that $\mathbb{E}[\boldsymbol{q}^{T}\boldsymbol{\epsilon}]=0$. We decompose $X$ as $X=U\Sigma V^{T}$ with $U\in\mathbb{R}^{d_{o}\times d_{a}}$ (which spans the column space of $X$), $\Sigma\in\mathbb{R}^{d_{a}\times d_{a}}$, and $V\in\mathbb{R}^{d_{a}\times d_{a}}$.
We will show that, when $\lambda\to+\infty$, $D=\left[\begin{matrix}U^{T}\\ 0\end{matrix}\right]$, B=[U,]B=[U,*], and $E=[V\Sigma^{-1},*]$ minimizes $\mathcal{L}_{a}$ where * indicates arbitrary entries and * vanishes when $d_{z}=d_{a}$.
First, this solution zeros the last two terms. For the last term, $D\boldsymbol{\epsilon}=\left[\begin{matrix}U^{T}\\ 0\end{matrix}\right]\boldsymbol{\epsilon}=0$ since $\boldsymbol{\epsilon}^{T}X=0$ and $U$ spans the column space of $X$. For the third term, $EDX=[V\Sigma^{-1},*]\left[\begin{matrix}U^{T}\\ 0\end{matrix}\right]U\Sigma V^{T}=I_{d_{a}}$. Then, since we have determined $D$, the second term becomes $\lVert(BD-I)\boldsymbol{\epsilon}\rVert_{2}^{2}=\lVert\boldsymbol{\epsilon}\rVert_{2}^{2}$. We can choose B=[U,]B=[U,*] to zero the first term since $BDX=[U,*]\left[\begin{matrix}U^{T}\\ 0\end{matrix}\right]X=X$. In this way, we obtain the minimal loss $\mathcal{L}^{*}_{a}=\lVert\boldsymbol{\epsilon}\rVert_{2}^{2}$. We can also verify that $B\boldsymbol{z}=B(C\boldsymbol{o}+D\boldsymbol{o}^{\prime})=BD\boldsymbol{q}=UU^{T}\boldsymbol{q}=\boldsymbol{q}$.